# Blameless Users in a Clean Room: Defining Copyright Protection for Generative Models

### v1

@@ 9 Discussion: provable copyright protection in practice?
People copy without generative models, and will continue copying with them. One way to evaluate the usefulness of a copyright-mitigation measure is to consider what happens when copying does occur. Who should pay? The existence of tainted NAF models muddies the makes it hard to assign culpability on the basis of NAF alone. DP gives a theoretical account of who is culpable: the provider if the data wasn’t golden; the user if not blameless (and possibly both or neither). But blamelessness may be impossible to check.
@@
Clean-room copy protection makes possible a simple indemnification rule: The model provider pays for infringement if the copied work appears too often in the training data. This is a question of fact that legal process and forensic experts are well-suited to resolve: experts for each party examine the training data and attempt to convince a fact finder. The rule gives users what they want. They get indemnity for copying when there is any possibility that the model is to blame, and can control their risk tolerance by controlling $\beta$. The rule gives model providers what they want too. Besides offering a valuable protection for users, the provider can trade-off their overall exposure to copyright penalties with their effort spent cleaning the data. Or the provider can pass the liability to third-party data providers contracted to provide golden data.
How reasonable are the requirements for clean room protection?
@@
Theorem 8.2 requires DP models, golden data, and blameless users. Are these requirements reasonable?
For pre-trained foundation models, DP and golden data are infeasible. DP generally requires more data for the same performance. And making golden dataset is much harder than deduplication (already a major challenge [LIN+21]): it depends on squishy legal standards and on data outside the dataset. Fine-tuning is more promising, because much less data is needed. The increased sample complexity of DP learning might be more manageable. And creating a golden dataset with tens of thousands of items seems feasible, possibly by commissioning new work.
@@
Blamelessness presents an even greater challenge: it is not checkable, not even by the user. Still, we believe that diligent users who are attentive to the possibility of inadvertent infringement can guarantee $\beta$-blamelessness for $\beta$ small enough to make Theorem 8.2 meaningful.
@@ Acknowledgements
We are indebted to Mayank Varia for many helpful discussions and encouragement. We thank Gautam Kamath, Yu-Xiang Wang, Seewong Oh, and especially Thomas Steinke for early discussions when the idea of this paper was still taking shape. Theorem 3.4 is based on an observation of Thomas. We thank Sarah Scheffler and the attendees of the 2024 Works-in-Progress Roundtable on Law and Computer Science at the University of Pennsylvania for feedback on an early draft.
@@ Appendix A Are we trying to reduce copyright to privacy? No!
Elkin-Koren, Hacohen, Livni, and Moran argue that copyright cannot be “reduced to privacy.” Referring to both NAF and DP by umbrella term algorithmic stability, they argue that the sort of provable guarantees that [VKB23] and this paper seek do not capture copyright’s complexities.
@@
We completely agree, and suspect that [VKB23] would too. The claim that “copyright can be reduced to privacy” is a straw man. It would be foolish to suggest that the whole of copyright law for generative AI be governed my a mathematical formalism like NAF or DP. That doesn’t mean that a mathematical formalism offering legal guarantees isn’t worthwhile—as a thought experiment or as a first step towards practical solutions.
@@ Appendix B Deferred material on near access-free models (Section 3)
This section presents our detailed treatment of near access-free models. It describes near access-freeness and its limitations. We prove that models can enable verbatim copying while still satisfying NAF. While NAF provides protection against a single prompt that is independent of the training data, it makes no guarantees against many prompts [LSK24], nor a single prompt derived from non-copyrightable ideas (not expression).
@@
B.1 Definitions and main results from [VKB23]
NAF is defined with respect to a function $\mathsf{safe}$. The $\mathsf{safe}$ function maps a copyrighted data point $c\in\mathcal{C}$ to a generative model $\mathsf{safe}_{c}\in\mathcal{P}$ trained without access to $c$. An example $\mathsf{safe}$ function is $\mathsf{sharded\mbox{-}safe}$. We simplify the notation of [VKB23] by fixing the divergence to maximum KL divergence.
@@ B.1 Definitions and main results from [VKB23]
Lemma B.3 ($k$-NAF event bound [VKB23]).
@@
As copying requires substantial similarity, bounding $k_{x}$ suffices to prevent violation whenever $\mathsf{safe}_{c}(\mathsf{SubSim}(c)|x)$ is negligibly small. Heuristically, we expect $\mathsf{safe}_{c}(\mathsf{SubSim}(c)|x)$ to be negligible in $|c|$. It may sometimes be large, e.g., when $x$ contains the copyrighted work $c$ [VKB23].
Theorem 3.3 presents the main feasibility result of [VKB23]: any training algorithm $\mathsf{Train}$ can be used as a black-box to construct an NAF model $\mathsf{CP}$, short for copy protection.
Theorem B.4 ([VKB23]).
@@
Algorithm 2 $\mathsf{sharded\mbox{-}safe}$ [VKB23]
@@
Algorithm 3 $\mathsf{CP}$: Copy-Protection [VKB23]
@@ B.2.1 $\mathsf{CP}$ can regurgitate training data
We show that $\mathsf{CP}$—the main NAF algorithm of [VKB23]—can fail to protect against copying. Using a prompt containing no copyrightable expression, a user can cause the NAF model returned by $\mathsf{CP}$ to regurgitate copyrighted training data. This is based on an observation of Thomas Steinke [Ste].
@@
Let $D$ be a dataset and $w\in D$. For ideas $\mathsf{id}\in\mathcal{I}$, let $D_{\mathsf{id}}=\{w^{\prime}\in D:\mathsf{ideas}(w^{\prime})=\mathsf{id}\}$. Let $\mathsf{Train}_{0}$ produce models with full support. Let $\mathsf{Train}_{\mathsf{kv}}$, $\mathsf{sharded\mbox{-}safe}_{\mathsf{kv}}$ and $\mathsf{CP}_{\mathsf{kv}}$ be as defined in Algorithms 1, 2, and 3, defined with respect to $\mathsf{Train}_{0}$ (and the preceding algorithms). Let $p\leftarrow\mathsf{CP}_{\mathsf{kv}}(D)$. Then
@@ B.2.2 $k$-NAF does not prevent full reconstruction
  • There exists a user $\mathsf{u}$ such that for all $D$ and $k>0$: $\mathsf{u}$ makes $\operatorname*{poly}(L,1/k)$ queries to $p_{k,D}$ and outputs $D$ with probability $>0.99$.
@@
We parse $D$ as a bit string of length $L$, and let $D[j]$ be its $j$th bit. For $k\geq 1$, $\mathsf{Train}$ outputs the model $p_{k,D}$ as follows:
$p_{k,D}(y|x)=\begin{cases}(D[x],D[x+1],\dots,D[x+\ell-1])&\text{if }x\in[L-\ell+1]\\ y\mathrel{{\leftarrow}\vbox{\hbox{\tiny\$}}}\{0,1\}^{\ell}&\text{otherwise}\end{cases}$
@@
For $10<k<$ 1, $\mathsf{Train}$ sets $\beta=2^{k}-1$ and outputs the model $p_{k,D}$ as follows:
$p_{k,D}(y|x)=\begin{cases}y\leftarrow\mathsf{Bernoulli}(\frac{1}{2}+\beta(D[x]-\frac{1}{2}))&\text{if }x\in[L]\\ y\leftarrow\mathsf{Bernoulli}(\frac{1}{2})&\text{otherwise}\end{cases}$
@@
To determine $D[j]$ with greater than $>1-\frac{1}{100L}$, the user $\mathsf{u}$ makes $\operatorname*{poly}(L,\log 1/\beta)=\operatorname*{poly}(L,1/k)$ queries to the model, and stores the majority. The user does this for each $j\in[L]$ and outputs the result. By a union bound, the user’s output is equal to $D$ with probability greater than $0.99$. ∎
B.3 VA3: an empirical attack on NAF [LSK24]
Li, Shen, and Kawaguchi propose and empirically evaluate a “Virtually Assured Amplification Attack” against NAF [LSK24]. At a high level, [LSK24] shows that an attacker, interacting with an NAF model in a black-box manner, can reliably induce a model to produce outputs that infringe on a target work $c^{*}$. They also give a white-box prompt writing algorithm for diffusion models (Anti-NAF) that greatly improves the performance of their attacks.
Overall, the work provides good evidence that the algorithms proposed by [VKB23] may not prevent infringement. Some uncertainty remains because the algorithm implemented in [LSK24] deviates from the original, as explained below. The paper leaves open whether $k$-NAF (with fixed $k$) prevents copyright infringement, formalizes a flawed attack model, and avoids the underlying definitional questions almost entirely.
We now describe [LSK24] in more detail, offering a somewhat different interpretation than the original.
@@ B.3 VA3: an empirical attack on NAF [LSK24]
More significant (for our purposes) is the existence of prompts $x$ whose one-shot probability of infringement is non-negligible: 0.40% or 8.52% in the previous example. The message is similar to our Theorem B.5. Namely, that the NAF constructions of [VKB23] admit models for which prompts $x$ derived from non-copyrightable aspects of $c^{*}$ can produce infringing generations. Based on the top half of [LSK24, Table 1], the prompts trivialize NAF’s protection by making $k_{x}\approx\log(1/\mathsf{safe}(\mathsf{SubSim}(c^{*})~{}|~{}x))$.
Though the paper does not speculate, we suspect that the mechanism for the failure is similar to our construction in Theorem B.5. The model, repeatedly fine-tuned on $c^{*}$, regurgitates $c^{*}$ with high enough probability that it survives the reweighting of $\mathsf{CP}$ and $\mathsf{CP\text{-}k}$ [LSK24, Fig. 8].
@@
First, the attack model has a conceptual flaw, highlighting the need for our definitions-first approach. The attacker knows $c^{*}$ and is actively trying to induce a similar generation [LSK24, Sec. 4.1]. Indeed, the Amplification Attack requires knowing $c^{*}$. This setup allows trivial attacks, like prompting $\mathtt{return\ c^{*}}$. Such an attack would not be meaningful. Any infringement would be the users’ fault, not something we care to prevent. Still, the flaw in the attack model doesn’t affect the paper’s experiments. Because they use a text-to-image model, he prompts used in the attack are just a few words long and contain nothing copyrightable.
Second, the results say nothing about $k$-NAF, where a fixed $k$ upper-bounds $k_{x}\leq k$ for every prompt $x$. At best, the experiments are negative results for the particular $k_{x}$-NAF algorithm $\mathsf{CP\text{-}k}$ given in [VKB23], where $k_{x}$ depends on $x$. But as we explain next, it is not entirely clear. (Our Theorems 3.4 and 3.5 are for $k_{x}$-NAF and $k$-NAF algorithms, respectively.)
Third, the algorithm in [LSK24] deviates from $\mathsf{CP\text{-}k}$ in an important way. As originally defined, $\mathsf{CP\text{-}k}$ is a rejection-sampling version of $\mathsf{CP}$ (Algorithm 3). Given a model $p$, a safe model $\mathsf{safe}$, and constant $k$, prompt $x$, and generation $y$, let $\rho(y|x)=\log\left(p(y|x)/\mathsf{safe}(y|x)\right)$. Oversimplifying, $\mathsf{CP\text{-}k}$ repeatedly samples $y\leftarrow p(\cdot|x)$ until $\rho(y|x)\leq k$, and returns the final sample. [VKB23] proves that $\mathsf{CP\text{-}k}$ is $k_{x}$-NAF for $k_{x}=k+\log(1/\nu(x))$, where $\nu(x)=\Pr_{y}[\rho(y|x)\leq k]$.
[LSK24]’s implementation differs. Instead of fixing $k$, they fix $\nu(x)$. The experiments sample many generations $y\leftarrow p(\cdot|x)$, and return the $\nu(x)=5\%,10\%,\ldots$ with smallest $\rho(y|x)$. This strikes us as a meaningful difference. At a minimum, the $\mathsf{CP\text{-}k}$ theorem doesn’t apply as is. Despite the gap, we view the results of [LSK24] as strong evidence of weaknesses of $\mathsf{CP\text{-}k}$. We conjecture that there is a value of $k_{x}$ for which the implemented version achieves $k_{x}$-NAF with high probability, but did not attempt to prove it.

### v2

@@ 9.1 Indemnification for genAI outputs
People copy without generative models, and will continue copying with them. One way to evaluate the usefulness of a copyright-mitigation measure is to consider what happens when copying does occur. Who should pay?
We are not asking where the liability would lie under existing law, but where it should lie. When generative AI is involved, holding a user strictly liable for all infringement is neither just nor effective. Punishment cannot incentivize users to avoid copying when avoiding copying is beyond their control.
It is hard to assign culpability on the basis of NAF alone, as tainted NAF models can induce a user acting appropriately to infringe. Clean-room copy protection gives a clearer theoretical account of who is culpable if training is differentially private. The provider is culpable if the data wasn’t golden; the user is culpable if not blameless; possibly both are culpable, or neither. However, this test is impossible to apply as blamelessness cannot in general be checked.
Even so, a simple indemnification rule is possible: The model provider pays for infringement if the copied work appears too often in the training data. This is a question of fact that legal process and forensic experts are well-suited to resolve. Experts for each party can examine the training data and testify before the finder of fact.
@@
Our indemnification rule gives users what they want. Users are indemnified when there is any possibility that the model is to blame, and they can control their risk tolerance by tuning $\beta$. The rule also gives model providers what they want. Besides offering a valuable protection for users, the provider can trade off their overall exposure to copyright penalties with their effort spent cleaning the data. Perhaps the provider can even pass the liability to third-party data providers contracted to provide golden data.
Many generative AI services already indemnify their users against copyright liability, with restrictions on user conduct. For example, OpenAI’s terms do not apply where the “Output was modified, transformed, or used in combination with products or services not provided by or on behalf of OpenAI” or where the user “did not have the right to use the Input or fine-tuning files to generate the allegedly infringing Output.” [^8] https://openai.com/policies/service-terms/ (accessed June 4, 2025).These are sensible restrictions. However, deciding whether the indemnification applies could require a complex forensic reconstruction and analysis of the creation of the infringing artifact, including how third-party tools were used and the role of the user’s prompts or fine-tuning. In contrast, our indemnification rule imposes no restrictions on the user and reduces the question of liability to a clear question of fact (albeit with the significant drawbacks of requiring golden datasets and differential privacy).
9.2 Is clean room protection practical?
@@
Theorem 8.2 requires DP models, golden data, and blameless users. Is it practical?
Making a golden dataset is much harder than deduplication (already a major challenge [lee2021deduplicating]): it depends on squishy legal standards and on data outside the dataset. Still, creating a golden dataset with tens of thousands of items seems feasible but expensive. One only needs to be able to determine if two items stem from a common in-copyright work (i.e., siblings in the copyright dependency graph). For a given pair of items, doing so with confidence is plausible with appropriate domain expertise. Another approach would be to create or commission new work with known copyright dependencies.
Constructing a golden dataset large enough to pre-train a foundation model is hopeless. Applications with more modest data needs are more promising, fine-tuning a pre-trained model for example. Even with differential privacy, tens of thousands of training data can achieve good utility.
@@
Blamelessness presents a greater challenge: it is not checkable, not even by the user. Still, we believe that diligent users who are attentive to the possibility of inadvertent infringement can guarantee $\beta$-blamelessness for $\beta$ small enough to make Theorem 8.2 meaningful.
@@ Acknowledgements
We are indebted to Mayank Varia for many helpful discussions and encouragement. We thank Gautam Kamath, Yu-Xiang Wang, Seewong Oh, and especially Thomas Steinke for early discussions when the idea of this paper was still taking shape. Theorem 3.4 is based on an observation of Thomas. We thank Sarah Scheffler, Randy Picker, Lior Strahilevitz, James Grimmelmann, anonymous reviewers, and attendees of the 2024 Works-in-Progress Roundtable on Law and Computer Science at the University of Pennsylvania for feedback on an early draft.
@@ Appendix A Are we trying to reduce copyright to privacy? No!
Elkin-Koren, Hacohen, Livni, and Moran argue that copyright cannot be “reduced to privacy.” Referring to both NAF and DP by umbrella term algorithmic stability, they argue that the sort of provable guarantees that [VyasKB23] and this paper seek do not capture copyright’s complexities.
@@
We completely agree, and suspect that [VyasKB23] would too. The claim that “copyright can be reduced to privacy” is a straw man. It would be foolish to suggest that the whole of copyright law for generative AI be governed my a mathematical formalism like NAF or DP. That doesn’t mean that a mathematical formalism offering legal guarantees isn’t worthwhile—as a thought experiment or as a first step towards practical solutions.
@@ Appendix B Deferred material on near access-free models (Section 3)
This section presents our detailed treatment of near access-free models. It describes near access-freeness and its limitations. We prove that models can enable verbatim copying while still satisfying NAF. While NAF provides protection against a single prompt that is independent of the training data, it makes no guarantees against many prompts [li2024va3], nor a single prompt derived from non-copyrightable ideas (not expression).
@@
B.1 Definitions and main results from [VyasKB23]
NAF is defined with respect to a function $\mathsf{safe}$. The $\mathsf{safe}$ function maps a copyrighted data point $c\in\mathcal{C}$ to a generative model $\mathsf{safe}_{c}\in\mathcal{P}$ trained without access to $c$. An example $\mathsf{safe}$ function is $\mathsf{sharded\mbox{-}safe}$. We simplify the notation of [VyasKB23] by fixing the divergence to maximum KL divergence.
@@ B.1 Definitions and main results from [VyasKB23]
Lemma B.3 ($k$-NAF event bound [VyasKB23]).
@@
As copying requires substantial similarity, bounding $k_{x}$ suffices to prevent violation whenever $\mathsf{safe}_{c}(\mathsf{SubSim}(c)|x)$ is negligibly small. Heuristically, we expect $\mathsf{safe}_{c}(\mathsf{SubSim}(c)|x)$ to be negligible in $|c|$. It may sometimes be large, e.g., when $x$ contains the copyrighted work $c$ [VyasKB23].
Theorem 3.3 presents the main feasibility result of [VyasKB23]: any training algorithm $\mathsf{Train}$ can be used as a black-box to construct an NAF model $\mathsf{CP}$, short for copy protection.
Theorem B.4 ([VyasKB23]).
@@
Algorithm 1 $\mathsf{sharded\mbox{-}safe}$ [VyasKB23]
@@
Algorithm 2 $\mathsf{CP}$: Copy-Protection [VyasKB23]
@@ B.2 Failures of NAF models
Parameters: Training algorithm $\mathsf{Train}_{0}$
Input: Data $D$
Output: Model $q_{\mathsf{kv}}$
1exLet $q_{0}\leftarrow\mathsf{Train}_{0}(D)$;
Initialize empty key-value store II, whose keys are ideas $\mathsf{id}$, values are sets of works $W\subset\mathcal{W}$;
for $w\in D$ do
   $I[\mathsf{ideas}(w)]\leftarrow I[\mathsf{ideas}(w)]\cup\{w\}$// add w to I[ideas(w)];
  1exLet $q_{\mathsf{kv}}$ the conditional generative model which on prompt $x$ does the following:;
   $W\leftarrow I[x]$;
   if $W\neq\emptyset$ then return $y$ sampled uniformly from $W$;
   else return $y$ sampled from $q_{0}(\cdot|x)$;
  1exResult: $q_{\mathsf{kv}}$
@@ B.2.1 $\mathsf{CP}$ can regurgitate training data
We show that $\mathsf{CP}$—the main NAF algorithm of [VyasKB23]—can fail to protect against copying. Using a prompt containing no copyrightable expression, a user can cause the NAF model returned by $\mathsf{CP}$ to regurgitate copyrighted training data. This is based on an observation of Thomas Steinke [steinkePersonal].
@@
Let $D$ be a dataset. For ideas $\mathsf{id}\in\mathcal{I}$, let $D_{\mathsf{id}}=\{w^{\prime}\in D:\mathsf{ideas}(w^{\prime})=\mathsf{id}\}$. Let $\mathsf{Train}_{0}$ produce models with full support. Let $\mathsf{Train}_{\mathsf{kv}}$, $\mathsf{sharded\mbox{-}safe}_{\mathsf{kv}}$ and $\mathsf{CP}_{\mathsf{kv}}$ be as defined in Algorithms 3, 1, and 2, defined with respect to $\mathsf{Train}_{0}$ (and the preceding algorithms). Let $p\leftarrow\mathsf{CP}_{\mathsf{kv}}(D)$. Then for all $w\in D$:
@@ B.2.2 $k$-NAF does not prevent full reconstruction
  • There exists a user $\mathsf{u}$ such that for all $D$ and $k>0$: $\mathsf{u}$ makes $\operatorname*{poly}(L,1/k)$ queries to $p_{k,D}$ and outputs $D$ with probability >0.99>0.99.
@@
We parse $D$ as a bit string of length $L$, and let $D[j]$ be its $j$th bit. We prove the result separately for $k\geq 1$ and $0<k<1$.
For $k\geq 1$, $\mathsf{Train}$ outputs the model $p_{k,D}$ as follows:
$p_{k,D}(\cdot|x)=\begin{cases}(D[x],D[x+1],\dots,D[x+\ell-1])&\text{if }x\in[L-\ell+1]\\ \text{sample uniform }y\in\{0,1\}^{\ell}&\text{otherwise}\end{cases}$
@@
For $0<k<1$, $\mathsf{Train}$ sets $\beta=2^{k}-1$ and outputs the model $p_{k,D}$ as follows:
$p_{k,D}(\cdot|x)=\begin{cases}y\sim\mathsf{Bernoulli}(\frac{1}{2}+\beta(D[x]-\frac{1}{2}))&\text{if }x\in[L]\\ y\sim\mathsf{Bernoulli}(\frac{1}{2})&\text{otherwise}\end{cases}$
@@
To determine $D[j]$ with greater than $1-\frac{1}{100L}$, the user $\mathsf{u}$ makes $\operatorname*{poly}(L,\log 1/\beta)=\operatorname*{poly}(L,1/k)$ queries to the model, and stores the majority. The user does this for each $j\in[L]$ and outputs the result. By a union bound, the user’s output is equal to $D$ with probability greater than $0.99$. ∎
B.3 VA3: an empirical attack on NAF [li2024va3]
Li, Shen, and Kawaguchi propose and empirically evaluate a “Virtually Assured Amplification Attack” against NAF [li2024va3]. At a high level, [li2024va3] shows that an attacker, interacting with an NAF model in a black-box manner, can reliably induce a model to produce outputs that infringe on a target work $c^{*}$. They also give a white-box prompt writing algorithm for diffusion models (Anti-NAF) that greatly improves the performance of their attacks.
Overall, the work provides good evidence that the algorithms proposed by [VyasKB23] may not prevent infringement. Some uncertainty remains because the algorithm implemented in [li2024va3] deviates from the original, as explained below. The paper leaves open whether $k$-NAF (with fixed $k$) prevents copyright infringement, formalizes a flawed attack model, and avoids the underlying definitional questions almost entirely.
We now describe [li2024va3] in more detail, offering a somewhat different interpretation than the original.
@@ B.3 VA3: an empirical attack on NAF [li2024va3]
More significant (for our purposes) is the existence of prompts $x$ whose one-shot probability of infringement is non-negligible: 0.40% or 8.52% in the previous example. The message is similar to our Theorem B.5. Namely, that the NAF constructions of [VyasKB23] admit models for which prompts $x$ derived from non-copyrightable aspects of $c^{*}$ can produce infringing generations. Based on the top half of [li2024va3, Table 1], the prompts trivialize NAF’s protection by making $k_{x}\approx\log(1/\mathsf{safe}(\mathsf{SubSim}(c^{*})~|~x))$.
Though the paper does not speculate, we suspect that the mechanism for the failure is similar to our construction in Theorem B.5. The model, repeatedly fine-tuned on $c^{*}$, regurgitates $c^{*}$ with high enough probability that it survives the reweighting of $\mathsf{CP}$ and $\mathsf{CP\text{-}k}$ [li2024va3, Fig. 8].
@@
First, the attack model has a conceptual flaw, highlighting the need for our definitions-first approach. The attacker knows $c^{*}$ and is actively trying to induce a similar generation [li2024va3, Sec. 4.1]. Indeed, the Amplification Attack requires knowing $c^{*}$. This setup allows trivial attacks, like prompting $\mathtt{return\ c^{*}}$. Such an attack would not be meaningful. Any infringement would be the users’ fault, not something we care to prevent. Still, the flaw in the attack model doesn’t affect the paper’s experiments. Because they use a text-to-image model, he prompts used in the attack are just a few words long and contain nothing copyrightable.
Second, the results say nothing about $k$-NAF, where a fixed $k$ upper-bounds $k_{x}\leq k$ for every prompt $x$. At best, the experiments are negative results for the particular $k_{x}$-NAF algorithm $\mathsf{CP\text{-}k}$ given in [VyasKB23], where $k_{x}$ depends on $x$. But as we explain next, it is not entirely clear. (Our Theorems 3.4 and 3.5 are for $k_{x}$-NAF and $k$-NAF algorithms, respectively.)
Third, the algorithm in [li2024va3] deviates from $\mathsf{CP\text{-}k}$ in an important way. As originally defined, $\mathsf{CP\text{-}k}$ is a rejection-sampling version of $\mathsf{CP}$ (Algorithm 2). Given a model $p$, a safe model $\mathsf{safe}$, and constant $k$, prompt $x$, and generation $y$, let $\rho(y|x)=\log\left(p(y|x)/\mathsf{safe}(y|x)\right)$. Oversimplifying, $\mathsf{CP\text{-}k}$ repeatedly samples $y\leftarrow p(\cdot|x)$ until $\rho(y|x)\leq k$, and returns the final sample. [VyasKB23] proves that $\mathsf{CP\text{-}k}$ is $k_{x}$-NAF for $k_{x}=k+\log(1/\nu(x))$, where $\nu(x)=\Pr_{y}[\rho(y|x)\leq k]$.
[li2024va3]’s implementation differs. Instead of fixing $k$, they fix $\nu(x)$. The experiments sample many generations $y\leftarrow p(\cdot|x)$, and return the $\nu(x)=5\%,10\%,\ldots$ with smallest $\rho(y|x)$. This strikes us as a meaningful difference. At a minimum, the $\mathsf{CP\text{-}k}$ theorem doesn’t apply as is. Despite the gap, we view the results of [li2024va3] as strong evidence of weaknesses of $\mathsf{CP\text{-}k}$. We conjecture that there is a value of $k_{x}$ for which the implemented version achieves $k_{x}$-NAF with high probability, but did not attempt to prove it.